e3:

The point-slope form of the equation of a line that passes through points (8, 4) and (0, 2) is y – 4 = y minus 4 equals StartFraction one-fourth EndFraction left-parenthesis x minus 9 right-parenthesis.(x – 8). What is the slope-intercept form of the equation for this line?

justus:

The slope-intercept form of the equation for this line is Y = 1/4x+ 6

e3:

The slope-intercept form of the equation of a line that passes through point (–2, –13) is y = 5x – 3. What is the point-slope form of the equation for this line?

justus:

The point-slope form of the equation for this line is y+13=5(x+2)

e3:

The ice skating rink charges an hourly fee for skating and $3 to rent skates for the day. Gillian rented skates and skated for 3 hours and was charged $21. Which equation represents the cost, c(x), of ice skating as a function of x, the number of hours of skating?

justus:

c(x) = 6x + 3